solve each system by substitution
4x-4y=-16
y=4x+22

Answer by Maths68(1474)

4x-4y=-16........(1)
y=4x+22 .........(2)
Put the value of y from (2) to (1)
4x-4y=-16
4x-4(4x+22)=-16
4x-16x-88=-16
4x-16x-88=-16
-12x=88-16
-12x=72
-12x/-12=72/-12
x=-6
Put the value of x in (2)
y=4(-6)+22 .........(2)
y=-24+22
y=-2
x,y=-6,-2

Check
=====
Put the value of x and y in (1)
4x-4y=-16
4(-6)-4(-2)=-16
-24+8=-16
-16=-16